metal has a threshold frequency of 5.85 x 1016 s-1, and after exposure to electromagnetic radiation an electron is emitted with a velocity of 6.95 x 105 m/s. What is the wavelength of the electromagnetic radiation used to eject the electron?  Mass of an electron is 9.109 x 10-31 kg
Answer #1

KE of electron = 0.5*m*v^2
= 0.5*(9.109*10^-31 Kg)*(6.95*10^5 m/s)^2
= 2.20*10^-19 J


Given:
fo = 5.85*10^16 s-1

use:
Eo = h*fo
=(6.626*10^-34 J.s)*(5.85*10^16) s-1
= 3.876*10^-17 J

UsE:
E = Eo + KE
= 3.876*10^-17 J + 2.20*10^-19 J
= 3.898*10^-17 J

Given:
E = 3.898*10^-17 J

use:
E = h*c/lambda
3.898*10^-17J =(6.626*10^-34 J.s)*(3.0*10^8 m/s)/lambda
lambda = 5.1*10^-9 m
Answer: 5.1*10^-9 m